3x=6x-15 then x+8=
Answer by tutorcecilia(2152)

3x=6x-15 [solve for the x-term=
3x-6x=6x-6x-15
-3x=-15
-3x/-3=-15/-3
x=5
[checking:
3(5)=6(5)-15
15=30-15
15=15 [checks out]
.
So, x=5 and
x+8= 5+8=13